Given that P(A) = 0.5, P(B) = 0.6, and P(A and B) = 0.30, determine P(A|B)
Answer by Jk22(389)

since P(A)P(B) = P(A&B) A and B are independent, hence the probability of A knowing B is the same as A, or

P(A|B) = P(A) = 0.5


Other way : P(A&B) = P(A|B)P(B), this implies : .3 = P(A|B)*0.6

-> P(A|B) = .3/.6 = 1/2 = 0.5